OpenStudy (rootbeer003):

Simplify square root 1/5

OpenStudy (anonymous):

\[\sqrt{\frac{1}{5}}\] like that?

OpenStudy (rootbeer003):

yuss

OpenStudy (rootbeer003):

will do

OpenStudy (rootbeer003):

i just like to check

OpenStudy (anonymous):

ok first off it is not "simplify" because there is no such thing, it is "write in simplest radical form" meaning get the radical out of the denominator \[\sqrt{\frac{1}{5}}=\frac{1}{\sqrt5}=\frac{1}{\sqrt5}\times \frac{\sqrt5}{\sqrt5}=\frac{\sqrt5}{5}\]

OpenStudy (rootbeer003):

oh

OpenStudy (rootbeer003):

harharhar

OpenStudy (anonymous):

the last number is the same as the first, the only difference is that the radical has moved from the denominator in to the numerator

OpenStudy (anonymous):

it is always true that \[\frac{1}{\sqrt{a}}=\frac{\sqrt{a}}{a}\]

OpenStudy (rootbeer003):

It wont let me message you btw

OpenStudy (anonymous):

so for example if you see \[\frac{1}{\sqrt2}\] you know it is \[\frac{\sqrt2}{2}\]

OpenStudy (rootbeer003):

also in the calculator, i got square root 5 0ver 5

OpenStudy (anonymous):

you got a hell of a calculator!!

OpenStudy (astrophysics):

Wow my calculator barely does fractions

OpenStudy (rootbeer003):

my sister has one from college

TheSmartOne (thesmartone):

ikr astro xD

OpenStudy (rootbeer003):

its like 90 bucks idk

OpenStudy (astrophysics):

The joke is I'm my own calculator! And I can barely do fractions, I'm dumb!

OpenStudy (anonymous):

oh look wolfram does the same thing http://www.wolframalpha.com/input/?i=1%2Fsqrt%282%29

OpenStudy (rootbeer003):

yaa mine looks like that

OpenStudy (rootbeer003):

but 5 over 5

OpenStudy (anonymous):

don't need to learn nutin honey, just plug it in to the wolf

OpenStudy (rootbeer003):

hahahaaaa

TheSmartOne (thesmartone):

LOL x'D

OpenStudy (rootbeer003):

ok lemme see if i have anymore annoying questions

OpenStudy (rootbeer003):

What is the x-intercept of the equation whose line is y = 3x – 12? –9 –4 4 9

OpenStudy (rootbeer003):

lol i cant

OpenStudy (anonymous):

what ?

OpenStudy (rootbeer003):

i eat too much

OpenStudy (anonymous):

x interecept, put \(y=0\) and solve \[3x-12=0\] in two steps

OpenStudy (rootbeer003):

4?

OpenStudy (anonymous):

yes 4

OpenStudy (rootbeer003):

What is the x-intercept of the equation whose line is 20x + 4y = 40? –10 –2 2 10

OpenStudy (rootbeer003):

10?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[20x=40\] solve for \(x\) here \(y=0\) like the last time

OpenStudy (rootbeer003):

2

OpenStudy (rootbeer003):

BINGO

OpenStudy (rootbeer003):

1.8

OpenStudy (rootbeer003):

IM SO SMART

OpenStudy (anonymous):

2, yes you are smart

OpenStudy (rootbeer003):

WHATS AN AMBASSADOR

OpenStudy (anonymous):

someone from one country who goes to another to drink wingspantails

TheSmartOne (thesmartone):

^ LOL but an openstudy ambassador is someone who welcomes new users on the site :)

OpenStudy (rootbeer003):

why do u make me read that

OpenStudy (rootbeer003):

OH

OpenStudy (anonymous):

wingspan tails lololz

OpenStudy (rootbeer003):

Solve 2x2 = 128. {±8} {±6} {±64} {±32}

OpenStudy (rootbeer003):

2x^2

OpenStudy (anonymous):

\[2x^2=128\] divide both sides by 2, what do you get as step one ?

OpenStudy (rootbeer003):

8?

OpenStudy (anonymous):

as a final answer, close \[x^2=64\\ x=\pm8\]

OpenStudy (rootbeer003):

so its 8?

TheSmartOne (thesmartone):

\(\bf plus~or~minus\) 8

OpenStudy (rootbeer003):

plus

OpenStudy (anonymous):

yeah your answer choices include the \(\pm\) so pick that one \(\pm8\)

OpenStudy (rootbeer003):

okie

OpenStudy (rootbeer003):

wolframalpha is a blessing

OpenStudy (anonymous):

wingspan tails, i love it going out to the bar to drink a few wingpan tails

OpenStudy (rootbeer003):

wanna know what i like

OpenStudy (rootbeer003):

taco sliders

OpenStudy (anonymous):

never heard of them sliders, yes taco sliders, no

OpenStudy (rootbeer003):

just sliders

OpenStudy (rootbeer003):

OK NEXT QUESTION

OpenStudy (anonymous):

...

OpenStudy (rootbeer003):

What number should be added to the expression x2 + 2x to change it into a perfect square trinomial? 1 2 3 4

OpenStudy (rootbeer003):

this one is a nono for me

OpenStudy (anonymous):

two steps half of two is ?

OpenStudy (rootbeer003):

1

OpenStudy (anonymous):

yes, and one squared is also one, so answer is one

OpenStudy (anonymous):

\[x^2+2x+1=(x+1)^2\]

OpenStudy (rootbeer003):

ooooooo k ur right

OpenStudy (rootbeer003):

finally

OpenStudy (anonymous):

i am usually wrong

OpenStudy (rootbeer003):

true true

TheSmartOne (thesmartone):

no he isn't :P